OpenStudy (anonymous):

Help me please! I have to Solve the Equation by Factoring: (3x-2)(4x+1)=0

OpenStudy (agreene):

(3x-2)=0 (4x+1)=0 3x=2 4x=-1 x= 2/3 x= -1/4 \[x=\frac23\space\space;\space\space x=-\frac14\]